40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15354     Entry time: Tue May 26 10:04:54 2020
Author: Jordan 
Type: Update 
Category: General 
Subject: N2 Replacement 

Replaced empty N2 tank, left tank at ~2000 psi, right tank ~2600 psi.

ELOG V3.1.3-